### v3

# How to Create a Flat Ten or Eleven Dimensional Space-time in the Interior of an Asymptotically Flat Four Dimensional String Theory

Abstract
By taking large mass and charge limit of a black hole in string theory we can create arbitrarily large regions where the space-time is approximately flat, but the moduli fields take values different from their asymptotic values. In this paper we describe a special case of this where black hole solutions in a four dimensional string theory, in the large mass and charge limit, can have an arbitrarily large region outside the horizon where a local observer will experience type IIA string theory in flat ten dimensional space-time. The curvature and other field strengths remain small everywhere between the asymptotic four dimensional observer and the ten dimensional region. By going to a different region of space, we can also get a large region where a local observer experiences M-theory in flat eleven dimensional space-time. By taking another solution in the same theory, one can create an arbitrarily large region where a local observer will experience type IIB string theory in flat ten dimensional space-time.
1 Introduction
In a previous paper[1] we had described a general procedure by which we can start with a string theory with a given set of values of the moduli fields at infinity and produce a background in which we have a large locally flat region where the moduli take different values. Thus an asymptotic observer can send out experimentalists to these regions to learn about the spectrum and S-matrix in a ‘different string theory’. This in fact demonstrates that the apparently different string theories characterized by different values of the moduli[2] are not different theories but different vacua of the same theory.
One example of this was discussed in [1] where it was shown how a large black hole in ten dimensional type IIA string theory carrying D0-brane charge allows us to create arbitrarily large, almost flat regions where the string coupling remains almost constant at any value between its asymptotic value and infinity. A large string coupling would correspond to M-theory in an almost flat eleven dimensional space-time[3, 4]. In this paper we describe another example where we begin with type IIA string theory compactified on $T^{6}$ and produce a background that contains a large, locally flat region where the size of the internal $T^{6}$ can be made arbitrarily large, keeping the string coupling fixed. Therefore a local observer in this region will experience type IIA string theory in ten dimensional flat space-time. There is another region in the same background where a local observer experiences M-theory in eleven dimensional flat space-time. The only price we pay is that larger the size of this ten or eleven dimensional region, the larger is the mass of the black hole needed to create such a configuration. Finally, by taking a different background in the same theory, one can construct an arbitrarily large region where a local observer experiences type IIB string theory in ten dimensional flat space-time.
The rest of the paper is organized as follows. In section 2 we describe a background where we have a locally flat region with large internal $T^{6}$, but the string coupling in that region is also large. This is not quite the configuration we want since we want the string coupling to remain small or finite. We remedy this problem in section 3 by placing the black hole solution of section 2 in the background of a second black hole. There we show that it is not only possible to identify a large region where the local physics is described by type IIA string theory in ten dimensional flat space-time, but it is also possible to identify another large region where the local physics is described by M-theory in eleven dimensional flat space-time. In section 4 we show, how by taking a different background in the same theory, we can produce a large region where a local observer experiences type IIB string theory in ten dimensional flat space-time with moderate or small string coupling. In section 5 we give an alternative construction by which we can achieve each of the three configurations – type IIA in ten dimensiona, M-theory in eleven dimensions and type IIB in ten dimensions by single black hole solutions. We end in section 6 with some comments.
2 D0-brane black hole in type IIA on $T^{6}$
We shall consider a four dimensional string theory obtained by compactifying type IIA string theory on $T^{6}$ and consider a non-rotating black hole in this theory carrying D0-brane charge. However in order to use the results of [5] directly, we shall first consider a black hole carrying D6-brane charge and then obtain the solution carrying D0-brane charge by T-dualizing all the compact directions. The relevant part of the ten dimensional action is given by:
$\int d^{10}x\sqrt{-\det G}\left[e^{-2\phi_{10}}\left\{R_{G}+4G^{\mu\nu}\partial_{\mu}\phi_{10}\partial_{\nu}\phi_{10}\right\}-G^{\mu\rho}G^{\nu\sigma}F_{\mu\nu}F_{\rho\sigma}\right]\,,$ (2.1)
where $F$ is the field strength associated with the RR one form gauge field, $G_{\mu\nu}$ is the string metric and $\phi_{10}$ is the dilaton field. The D6-brane is magnetically charged under the RR one form gauge field, and the black hole solution carrying such charges is given by[5]:
$\displaystyle F$ $=$ $Q\,e^{-\phi_{0}}\,\epsilon_{2}$
$ds^{2}$ $=$ $-\left[1-\left({r_{+}\over r}\right)\right]\,\left[1-\left({r_{-}\over r}\right)\right]^{\gamma_{x}-1}dt^{2}+\left[1-\left({r_{+}\over r}\right)\right]^{-1}\,\left[1-\left({r_{-}\over r}\right)\right]^{\gamma_{r}}dr^{2}$
$+$ $r^{2}\,\left[1-\left({r_{-}\over r}\right)\right]^{\gamma_{r}+1}d\Omega_{2}^{2}+\left[1-\left({r_{-}\over r}\right)\right]^{\gamma_{x}}\,\sum_{i=4}^{9}dx^{i}dx^{i}\,,$
$e^{-2\phi_{10}}$ $=$ $e^{-2\phi_{0}}\,\left[1-\left({r_{-}\over r}\right)\right]^{\gamma_{\phi}}\,,\qquad r_{-}\leq r_{+}\,,$ (2.2)
where $\epsilon_{2}$ is the volume form on a unit two sphere whose metric is denoted by $d\Omega_{2}^{2}$ and [^1] The case we are considering corresponds to the $\alpha=0$, $D=4$ case of [5].
$\gamma_{x}={1\over 2}\,,\qquad\gamma_{r}={1\over 2}\,,\qquad\gamma_{\phi}=-{3\over 2}\,.$ (2.3)
$x^{4},\cdots,x^{9}$ are coordinates along the compact directions, which we shall take to be periodic with period $2\pi$. The $e^{-\phi_{0}}$ factors were not present in the solution given in [5], but we have included it here using the scaling ‘symmetry’ $\phi_{10}\to\phi_{10}+\phi_{0}$, $F\to e^{-\phi_{0}}F$ under which the action scales by $e^{-2\phi_{0}}$ and the equations of motion remain unchanged. The parameters $Q$ and the mass $M$ of the black hole solution are related to $r_{\pm}$ via the relations[5]:
$Q={1\over 2}(r_{+}r_{-})^{1/2},\qquad M=r_{+}-{1\over 2}r_{-}\,.$ (2.4)
The horizon of this black hole is at $r=r_{+}$. Extremal limit would correspond to $r_{-}\to r_{+}$, but for now we keep $r_{-}$ and $r_{+}$ as independent parameters.
Now any two derivative action in $d$ dimensions containing metric $G_{\mu\nu}$, and $p$-forms $A^{(p)}$ for different values of $p$, scales by a factor of $\lambda^{d-2}$ under the scaling
$G_{\mu\nu}\to\lambda^{2}\,G_{\mu\nu},\qquad A^{(p)}\to\lambda^{p}\,A^{(p)}\,.$ (2.5)
This takes a solution to a new solution. In the new solution any invariant constructed from these fields with two derivatives, e.g. the curvature scalar or $G^{\mu\rho}G^{\nu\sigma}F_{\mu\nu}F_{\rho\sigma}$ will scale as $\lambda^{-2}$. Invariants containing larger number of derivatives will scale as larger powers of $\lambda^{-1}$. Thus in the large $\lambda$ limit all these invariants will approach zero and the space-time will appear to be locally flat. We shall carry out this rescaling for the solution (2) by regarding this as a solution in four dimensional theory. This means that we do not scale the internal components of the metric, rather we regard the sizes of the internal circles as scalar moduli fields in the four dimensional theory and hold them fixed as we scale the other components of the metric by $\lambda^{2}$. In the solution (2) this will correspond to scaling $r$, $t$, $r_{+}$ and $r_{-}$ by $\lambda$ [^2] While the scaling of $r$ and $t$ are just coordinate transformations, the scaling of $r_{\pm}$ generate genuinely different solution. We shall not display $\lambda$ explicitly but keep in mind that we work in the region where $r_{\pm}$ are large and $r$ is of order $r_{+}$. It is easy to check that in this case the metric is locally approximately flat and the magnitude of the field strength is approximately zero. Total integrated flux however scales as $\lambda$ and is large.
We now define
$a\equiv r_{+}-r_{-},\qquad b=r-r_{+}\,,$ (2.6)
and take $a/r_{+}$ and $b/r_{+}$ to be small but fixed as we scale $r_{\pm}$ to large values. This means that we stay in the near horizon region of a large, near extremal black hole. The local geometry is still nearly flat as long as we take $r_{+}$ to be sufficiently large.
We now compute the scalar $\phi_{10}$ and the radii $R$ of the compact directions in this region. They are given by
$e^{-2\phi_{10}}=e^{-2\phi_{0}}\left({r_{+}+b\over b+a}\right)^{3/2}\,,\qquad R=\left({b+a\over r_{+}+b}\right)^{1/4}\,.$ (2.7)
In particular the four dimensional dilaton $\phi_{4}$, is given by
$e^{-2\phi_{4}}\equiv e^{-2\phi_{10}}\,R^{6}=e^{-2\phi_{0}}\,.$ (2.8)
Thus $\phi_{4}$ does not flow and its value at $r=r_{+}+b$ is the same as at infinity.
From (2.7) we see that for $b,a<<r_{+}$, the sizes of the compact directions, measured in the string metric, is small and $e^{-2\phi_{10}}$ is large, which means that the string coupling is small. We shall now go to the T-dual description of this geometry, by making an $R\to\widetilde{R}=1/R$ transformation on all the circles. During this transformation the four dimensional dilaton $\phi_{4}$ remains the same, $\phi_{10}$ transforms to $\widetilde{\phi}_{10}$ accordingly, and the integrated flux of gauge field is reinterpreted as the D0-brane charge of the black hole. The near horizon moduli now take the form: [^3] Note that we could get this result directly by solving the equations of motion of low energy supergravity. The only reason for going via D6-branes is to lift the solution from [5].
$\widetilde{R}=R^{-1}=\left({r_{+}+b\over b+a}\right)^{1/4},\qquad e^{2\widetilde{\phi}_{10}}=e^{2\phi_{4}}\,\widetilde{R}^{6}=e^{2\phi_{0}}\,\left({r_{+}+b\over b+a}\right)^{3/2}\,.$ (2.9)
We also note that in the new duality frame, the asymptotic radii of the circles remain 1 and the dilaton $\widetilde{\phi}_{10}$ takes value $\phi_{0}$.
We see from (2.9) that for $b,a<<r_{+}$, the radii $\widetilde{R}$ of the compact directions, measured in the string metric, is large. However the string coupling $e^{\widetilde{\phi}_{10}}$ in this region is large. Hence we still have not reached our goal, which is to produce a region of space-time where the internal radii can be made (arbitrarily) large and the string coupling can be kept small or finite.
To address this issue, we note that if we are allowed to adjust $\phi_{0}$, then we could take $e^{2\phi_{0}}$ to be small to compensate for the large contribution from the last factor in the expression for $e^{2\widetilde{\phi}_{10}}$ in (2.9). However we would like to get a large ten dimensional space-time in a given four dimensional string theory for which all the asymptotic moduli are fixed, including the coupling $e^{\phi_{0}}$. So $\phi_{0}$ cannot be adjusted to achieve our goal. In the next section we shall describe a different black hole solution for which we can identify arbitrarily large locally flat regions of space-time where the string coupling can be made arbitrarily small, keeping all other moduli finite. Thus by placing the black hole considered in this section (the first black hole) in the background of the second black hole that we shall describe in the next section, we can vary the effective $\phi_{0}$ that enters the computation of this section by varying the position of the first black hole. This way we’ll be able to achieve our goal of creating a region where the space-time is approximately ten dimensional and have finite ten dimensional string coupling.
3 Black hole with winding and momentum charge
In this section we shall analyze a non-rotating black hole solution in type IIA string theory on $T^{6}$, carrying fundamental string winding and momentum charge along one of the circles. For simplicity we shall take the momentum and winding charges to be equal. We also take the asymptotic radii of the circles to be 1 and the asymptotic string coupling to be 1. The solution is obtained from the solutions given in [6] and reviewed in appendix A by setting $\alpha=0$ and taking the unit vector $\vec{p}$ to be along the circle carrying the winding and momentum charges. The four dimensional string metric $ds_{4}^{2}$, the ten dimensional dilaton $\phi_{10}$, the four dimensional dilaton $\phi_{4}$, the $12\times 12$ symmetric, $SO(6,6)$ matrix valued scalar $M$ and the non-vanishing component of the gauge fields are given by:
$\displaystyle ds_{4}^{2}$ $=$ $-\Delta^{-1}\rho^{2}(\rho^{2}-2m\rho)dt^{2}+\rho^{2}(\rho^{2}-2m\rho)^{-1}d\rho^{2}+\rho^{2}d\Omega_{2}^{2}\,,$
$e^{-2\phi_{10}}$ $=$ $e^{-2\phi_{4}}=\Delta^{1/2}\,\rho^{-2}\,,$
$M$ $=$ $I_{12}\,,$
$A_{t}$ $=$ $-{1\over\sqrt{2}}\,m\,\rho\,\sinh\beta\,\Delta^{-1/2}\,,$
$\Delta$$\equiv$ $(\rho^{2}+m\rho(\cosh\beta-1))^{2}\,.$ (3.1)
The horizon is at $\rho=2m$. The $M=I_{12}$ result, $I_{12}$ being the $12\times 12$ identity matrix, implies that the radii of the compact directions remain constant at 1 for all $\rho$. $A_{t}$ is the gauge field that couples to the momentum plus winding charge. The extremal limit corresponds to $m\to 0$, $\beta\to\infty$ with $m\sinh\beta$ fixed. We shall not take this limit for now.
The scaling that makes the metric almost locally flat and the invariant tensors almost zero is $m\to\lambda^{\prime}m$, $\rho\to\lambda^{\prime}\rho$, $t\to\lambda^{\prime}t$ for large $\lambda^{\prime}$. Again we shall not display the factors of $\lambda^{\prime}$ explicitly but keep in mind that we work with large $m$ and $\rho$ with $\rho/m>2$ fixed. Also we shall take $\beta$ large but fixed. In this limit the geometry becomes nearly flat and all the invariant tensors constructed from two or more derivatives of fields become nearly zero. The dilaton $\phi_{10}$ has the form:
$e^{-2\phi_{10}}\simeq m\,\rho^{-1}\cosh\beta\,.$ (3.2)
We now see that for large $\beta$ with fixed $\rho/m>2$, $e^{-2\phi_{10}}$ is large and hence $\phi_{10}$ becomes large and negative. Therefore in this region the string coupling $e^{\phi_{10}}$ is small. If we take the parameter $\lambda^{\prime}$ to be much larger than the parameter $\lambda$ used in the scaling in section 2, then the size of the black hole described in section 2, even though large, is small compared to the distance scale over which the fields change appreciably in the black hole background given in (3). Thus the black hole of section 2 will appear to be a point object from the perspective of the background 3 and it makes sense to place it at any point in the background (3). In particular we can place it at $\rho\sim m$. This configuration will evolve with time, but the time scale over which the fields will evolve will be of order $\lambda^{\prime}$ that can be made arbitrarily large. In that case $\phi_{0}$ appearing in (2.9) will be given by the $\phi_{10}$ in (3.2). Substituting this into (2.9) gives
$\widetilde{R}=\left({r_{+}+b\over b+a}\right)^{1/4},\qquad e^{2\widetilde{\phi}_{10}}={\rho\over m\cosh\beta}\,\,\left({r_{+}+b\over b+a}\right)^{3/2}\,.$ (3.3)
Taking $a,b<<r_{+}$, $\rho\sim m$ and $\beta\sim\cosh^{-1}[\{r_{+}/(b+a)\}^{3/2}]>>1$, we can achieve our goal of getting large $\widetilde{R}$ and finite $\widetilde{\phi}_{10}$ i.e. a locally flat geometry where space-time is ten dimensional and the string coupling is finite.
It is also interesting to explore the parameters that arise if we regard this as an eleven dimensional $M$-theory compactification using the duality between type IIA string theory in ten dimensions and circle compactification of $M$-theory[3, 4]. Let $R_{M}$ denote the radii of the $x^{i}$ directions for $4\leq i\leq 9$ in the eleven dimensional metric and $R_{11}$ denote the radius of the emergent circle as we go from type IIA to M-theory description. In this case, we have[4]
$R_{11}=e^{2\widetilde{\phi}_{10}/3}=\left({\rho\over m\cosh\beta}\right)^{1/3}\left({r_{+}+b\over b+a}\right)^{1/2}\,,\qquad R_{M}=e^{-\widetilde{\phi}_{10}/3}\,\widetilde{R}=\left({m\cosh\beta\over\rho}\right)^{1/6}\,.$ (3.4)
From this we see that by taking $\beta$ and $r_{+}/(a+b)$ large, and appropriately choosing their ratios, it is possible to make both $R_{11}$ and $R_{M}$ large. In that case we would create a large region where a local observer would experience eleven dimensional M-theory in almost flat space-time. In fact, for the same geometry, characterized by large values of $r_{+}/a$ and $\beta$, we can vary $b\equiv r-r_{+}$, labelling the region, to have either $\widetilde{R}$ large with $\widetilde{\phi}_{10}$ small or finite (IIA in flat space-time), or $R_{11}$ and $R_{M}$ large (M-theory in flat space-time).
Since the constructions described above involve taking various combinations large, we shall now summarize the order of limits needed for getting the ten and eleven dimensional regions:
  1. 1.
    Type IIA in flat ten dimensional space-time:
    $m>>r_{+}>>>r_{+}/(b+a)\sim(\cosh\beta)^{2/3}>>1,\qquad\rho-2m\sim m\,.$ (3.5)
    We remind the reader that $a/r_{+}\equiv(r_{+}-r_{-})/r_{+}$ is the non-extremality parameter of the first black hole, $1/\cosh\beta$ is the non-extremality parameter of the second black hole, $r=r_{+}+b$ labels the region around the first black hole where we work and $\rho$ labels the region around the second black hole where we work. Also we have set $\hbar=1$, $c=1$, $\alpha^{\prime}=1$. The $m>>r_{+}$ condition ensures that the second black hole is much larger than the first black hole. The $m,r_{+}>>>r_{+}/(a+b),(\cosh\beta)^{2/3}$ condition ensures that the curvature and other invariants from the perspective of four dimensional space-time are made small enough so that even if the natural invariants from the higher dimensional perspective carry some powers of the sizes of the extra dimensions, these invariants still remain small. So these inequalities should really be interpreted as taking $m$ and $r_{+}$ to be larger than any power of $r_{+}/(a+b)$ and $(\cosh\beta)^{2/3}$ that might arise in the expressions for ten dimensional invariants constructed from derivatives of the metric and other fields. This is what is meant by the $>$much-greater-thanabsent>>>> > > symbol.
  2. 2.
    M-theory in flat eleven dimensional space-time:
    $m>>r_{+}>>>r_{+}/(b+a)>>(\cosh\beta)^{2/3}>>1,\qquad\rho-2m\sim m\,.$ (3.6)
    As in the ten dimensional case, we need to take $m$ and $r_{+}$ to be larger than any power of $r_{+}/(a+b)$ and $(\cosh\beta)^{2/3}$ that might arise in the expression for the eleven dimensional invariants constructed from derivatives of metric and other fields.
4 Black hole with purely winding charge
In this section we shall describe how to construct a configuration that contains an arbitrarily large region where the local observer experiences type IIB string theory in flat ten dimensional space-time. For this we need to make use of a third black hole solution in type IIA string theory on $T^{6}$ that carries only winding charge along one of the circles. We shall take this to be the ninth circle. This solution can again be read out from the general solution given in appendix A by setting $\alpha=\beta$ and taking the vectors $\vec{n}$ and $\vec{p}$ to be directed along the same compact direction $x^{9}$. Using (A)-(A.3), the four dimensional dilaton $\phi_{4}$ and the radius $R_{9}$ of the ninth direction takes the form:
$e^{2\phi_{4}}=e^{2\widehat{\phi}_{0}}\,(1+2\widehat{m}\widehat{\rho}^{-1}\sinh^{2}\alpha)^{-1/2}\,,$ (4.1)
$R_{9}=(1+2\widehat{m}\widehat{\rho}^{-1}\sinh^{2}\alpha)^{-1/2}\,,$ (4.2)
where we have introduced new parameter $\widehat{m}$ and new radial variable $\widehat{\rho}$ to distinguish them from the variables used in section 3. The other radii are kept fixed at 1. ${\widehat{\phi}_{0}}$ denotes the asymptotic value of $\phi_{4}$ for this solution. This is also the asymptotic value of the ten dimensional dilaton $\phi_{10}$ for this solution since $R_{9}$ approaches 1 in this limit and the other radii are also fixed at 1. On the other hand at a generic position labelled by $\widehat{\rho}$, the ten dimensional dilaton $\phi_{10}$ takes the form
$e^{2\phi_{10}}=e^{2\phi_{4}}\,R_{9}=e^{2\widehat{\phi}_{0}}\,(1+2\widehat{m}\widehat{\rho}^{-1}\sinh^{2}\alpha)^{-1}\,.$ (4.3)
The horizon of this solution is at $\widehat{\rho}=2\widehat{m}$. Hence we shall stay in the region $\widehat{\rho}>2\widehat{m}$.
Let us now put this solution in the background of the solution given in (3) at the position $\rho$, taking the size of (3) to be much larger than the size of this solution (i.e. we take $m,\rho>>\widehat{m},\widehat{\rho}>>1$). Then we can use (3.2) to write
$e^{-2\widehat{\phi}_{0}}=m\,\rho^{-1}\,\cosh\beta\,.$ (4.4)
Taking $\widehat{m}\widehat{\rho}^{-1}\sinh^{2}\alpha$ to be large, we get, from (4.3), (4.2):
$e^{2\phi_{10}}={1\over 2}\,m^{-1}\,\rho\,(\cosh\beta)^{-1}\,\widehat{m}^{-1}\,\widehat{\rho}\,(\sinh\alpha)^{-2}\,,\qquad R_{9}={1\over\sqrt{2}}\,\widehat{m}^{-1/2}\,\widehat{\rho}^{1/2}\,(\sinh\alpha)^{-1}\,.$ (4.5)
We now put the first black hole, described in section 2, at the position $\widehat{\rho}$ by taking $m,\rho>>\hat{m},\hat{\rho}>>r_{\pm},a,b>>1$. This amounts to replacing $e^{2\phi_{0}}$ in (2.9) by $e^{2\phi_{10}}$ given in (4.5). Also, the radius of the ninth circle will now be multiplied by $R_{9}$ given in (4.5) since this becomes the asymptotic value of this radius from the perspective of the solution (2.9). This gives the following solution:
$\widetilde{R}=\left({r_{+}+b\over b+a}\right)^{1/4},\qquad\widetilde{R}_{9}=\left({r_{+}+b\over b+a}\right)^{1/4}{1\over\sqrt{2}}\,\widehat{m}^{-1/2}\,\widehat{\rho}^{1/2}\,(\sinh\alpha)^{-1},$
$e^{2\widetilde{\phi}_{10}}={1\over 2}\,m^{-1}\,\rho\,(\cosh\beta)^{-1}\,\widehat{m}^{-1}\,\widehat{\rho}\,(\sinh\alpha)^{-2}\,\left({r_{+}+b\over b+a}\right)^{3/2}\,.$ (4.6)
Note that if we ignore the ${1\over 2}\,\widehat{m}^{-1}\,\widehat{\rho}\,(\sinh\alpha)^{-2}$ factor, then the solution reduces to (3.3).
The final step is to make a $R\to R^{-1}$ duality transformation on the ninth circle to map this to a type IIB string theory. Denoting the new variables by bars, we get
$\overline{R}=\widetilde{R}=\left({r_{+}+b\over b+a}\right)^{1/4},\qquad\overline{R}_{9}=\widetilde{R}_{9}^{-1}=\left({r_{+}+b\over b+a}\right)^{-1/4}{\sqrt{2}}\,\widehat{m}^{1/2}\,\widehat{\rho}^{-1/2}\,\sinh\alpha\,,$
$e^{2\overline{\phi}_{10}}=e^{2\widetilde{\phi}_{10}}\widetilde{R}_{9}^{-2}=m^{-1}\,\rho\,(\cosh\beta)^{-1}\,\left({r_{+}+b\over b+a}\right)\,.$ (4.7)
In order to get type IIB string theory with large compactification radii and weak or moderate coupling, we need $\overline{R},\overline{R}_{9}>>1$ and $e^{\overline{\phi}_{10}}\stackrel{{\scriptstyle<}}{{\sim}}1$. If we choose $\rho\sim m$ and $\widehat{\rho}\sim\widehat{m}$, then this requires:
$r_{+}>>a,b,\qquad\sinh\alpha>>\left({r_{+}+b\over b+a}\right)^{1/4},\qquad\cosh\beta\stackrel{{\scriptstyle>}}{{\sim}}\left({r_{+}+b\over b+a}\right)\,.$ (4.8)
The analog of (3.5) now takes the form:
$m>>\widehat{m}>>r_{+}>>>\sinh^{4}\alpha>>r_{+}/(b+a)\sim\cosh\beta>>1,\qquad\rho\sim m,\qquad\widehat{\rho}\sim\widehat{m}\,.$ (4.9)
Finally, we note that if we had used the type IIB description from the begining, then the first black hole would carry D1-brane charge along the ninth direction, the second black hole would carry equal amount of fundamental string winding and momentum charges along one of the circles and the third black hole would carry momentum along the ninth direction.
5 Achieving the goals with single black holes
In the constructions described above we made use of multiple black holes to produce arbitrarily large regions where a local observer experiences ten dimensional type IIA or type IIB string theories or eleven dimensional M-theory. It is natural to ask if this can be achieved using a single black hole. In this section we can describe how this can be done.
For this analysis we shall make use of black holes carrying charges representing momenta along the compact directions. We shall first describe the construction in type IIA string theory on $T^{6}$. This can again be obtained as a special case of the solution given in appendix A where we set $\alpha=-\beta$ and take the vectors $\vec{p}=\vec{n}$ to be some generic six dimensional unit vector. Then the four dimensional metric $ds_{4}^{2}$, the four dimensional dilaton $\phi_{4}$ and the internal six dimensional metric $ds_{6}^{2}$ describing the moduli fields have the following functional dependence on the radial variable $\rho$::
$\displaystyle ds_{4}^{2}$ $=$ $-{\rho-2m\over\rho+2m\sinh^{2}\alpha}dt^{2}+{\rho\over\rho-2m}d\rho^{2}+\rho^{2}d\Omega_{2}^{2}\,,$
$e^{2\phi_{4}}$ $=$ $e^{2\phi_{0}}\,(1+2m\rho^{-1}\sinh^{2}\alpha)^{-1/2},$
[…]
where $\phi_{0}$ is the asymptotic value of the dilaton $\phi_{4}$. The horizon of the solution is at $\rho=2m$ and the momentum charge vector carried by the black hole is proportional to $m\,\sinh^{2}\alpha\,\vec{n}$. As before, we shall take the large $m$ limit at fixed $\rho/m$ so as to smoothen the geometry and make it locally flat. From (5) we can also compute the ten dimensional dilaton
$e^{2\phi_{10}}=e^{2\phi_{4}}\,R=e^{2\phi_{0}}\,.$ (5.2)
Thus the ten dimensional dilaton is constant. This shows that this is a purely gravitational solution and hence can be easily lifted to any theory, e.g. M-theory on $T^{7}$ or type IIB on $T^{6}$.
By taking $\rho\sim m$ and $\alpha$ large, we can take $R$ to be large. In this case, apparently (5) indicates that only one of the directions $\vec{n}$ of the torus becomes large and the orthogonal directions remain finite. For example, if each $y^{i}$ is periodic with period one, and if $\vec{n}\propto(1,1,\cdots 1)$, then the cycle that shifts $y^{1}$ by $2\pi$ and $y^{2}$ by $-2\pi$ will have finite size since it will not be affected by the term proportional to $R^{2}$ in the metric. However by choosing $\vec{n}$ appropriately we can ensure that all closed cycles on the torus have large size. We shall illustrate this by an example below. Let us take
$\vec{n}=(1,c,c^{2},\cdots c^{d-1})/\sqrt{1+c^{2}+\cdots+c^{2d-2}}$ (5.3)
[…]
$2\pi\left[m_{1}^{2}+\cdots+m_{d}^{2}+(R^{2}-1)(m_{1}+m_{2}c+\cdots m_{d}c^{d-1})^{2}/(1+c^{2}+\cdots+c^{2d-2})\right]^{1/2}\,.$ (5.4)
For generic $m_{i}$ the cycle will have size of order $R$. We could avoid it by choosing the $m_{i}$’s so that $(m_{1}+m_{2}c+\cdots m_{d}c^{d-1})$ is small. This can be done in two ways: either set some of the $m_{i}$’s to zero or choose the $m_{i}$’s such that the different terms in the sum cancel. In the first approach the smallest value is achieved by setting $m_{i}=0$ for $i\neq d-1$ and $m_{d}=1$. This gives a cycle size of order $2\pi Rc^{(d-1)}\sim R^{1/d}$ for $c\sim R^{-1/d}$. Thus we have a large size for large $R$. On the other hand in order to have cancellation between different terms in $(m_{1}+m_{2}c+\cdots m_{d}c^{d-1})$, some of the $m_{i}$’s must be of order $c^{-1}$. Then the $R$ independent terms in (5.4) will have large contribution and we still will have a minimum cycle size of order $R^{1/d}$.
One can also see this by analyzing the spectrum of momentum carrying states. If $\vec{k}$ denotes the quantized momentum along $T^{d}$ taking value on a lattice of integers, then for the metric given in (5), its contribution to the mass2 in four space-time dimensions is given by:
$m^{2}=G^{ij}k_{i}k_{j}=\vec{k}^{2}+(R^{-2}-1)(\vec{n}.\vec{k})^{2}\,.$ (5.5)
Let us for definiteness take
$R=N^{d},\qquad c=N^{-1}\,,$ (5.6)
for some large integer $N$. This can be achieved by taking the momenta along the $d$ circles to be some multiple of $(1,N,N^{2},\cdots N^{d-1})$. We now introduce the following basis states in the momentum lattice:
$\vec{e}_{1}=(1,0,\cdots 0),\quad\vec{e}_{2}=(N,1,0,\cdots 0),\quad\vec{e}_{3}=(N^{2},N,1,0,\cdots 0),\quad\cdots,\ \vec{e}_{d}=(N^{d-1},N^{d-2},\cdots 1)\,.$ (5.7)
We also have from (5.3)
$\vec{n}=(1+N^{2}+\cdots N^{2d-2})^{-1/2}(N^{d-1},N^{d-2},\cdots 1)\,.$ (5.8)
Then a general momentum vector may be expressed as
$\vec{k}=\sum_{i=1}^{d}q_{i}\vec{e}_{i},\qquad q_{i}\in{\hbox{ Z\kern-4.55244ptZ}}\,.$ (5.9)
Substituting this into (5.5), we get
$m^{2}=N^{-2}\sum_{i=1}^{d}q_{i}^{2}+{\cal O}(N^{-3})\,.$ (5.10)
This agrees with the spectrum of a theory where each of the $d$ circles is compactified on a circle of radius $N=c^{-1}$.
This achieves the goal of having large dimensions of the torus. This can be applied to each of the three theories. In particular, by regarding the original four dimensional theory as type IIA on $T^{6}$ and considering a black hole that carries large six dimensional momentum charge we can get type IIA in flat ten dimensional space-time. The same theory can be regarded as M-theory on $T^{7}$ and now by switching on appropriate large seven dimensional momentum charge (which corresponds to six dimensional momentum charge and D0-brane charge in the original type IIA description) we can get M-theory in flat eleven dimensional space-time. The same theory can also be regarded as type IIB on $T^{6}$ via a T-duality transformation and by switching on large six dimensional momentum charge (which corresponds to momentum along five of the circles and fundamental string winding along the sixth circle in the type IIA description) we can get type IIB theory in flat ten dimensional space-time.
6 Discussions
In [1] we had described a general strategy for producing almost locally flat background where the moduli fields take a different value compared to the values they take at infinity. In this paper we have described a particular example of this mechanism where starting from a four dimensional string theory with finite values of all the moduli fields, we have described backgrounds that have large locally flat regions where the space-time is effectively ten dimensional and the string coupling remains finite and another background that has a large locally flat region where the space-time is effectively eleven dimensional.
This of course does not demonstrate that we can reach all possible values of the moduli this way. It will be interesting to try to establish this or find the limitations if any. In particular the moduli space of string theory has interesting topology changing transitions and one would like to realize them as part of space-time background. One interesting case of these is the conifold transition[7, 8] – that takes a deformed conifold to a resolved conifold. One could ask if it is possible to construct a background where in some region of space-time the internal manifold is a resolved conifold and in another region it is a deformed conifold. Since this involves both vector multiplet and hypermultiplet moduli of $N=2$ supersymmetric string theory, the configuration is likely to involve both black holes and loops of black strings.
Finally we note that while the asymptotic geometry is taken to be flat space-time (times an internal manifold) the geometries that we can construct in the interior are not limited to flat space-time. For example in ten dimensional type IIB string theory we can take a stack of $N$ spherical D3-branes so that it has finite energy, but take the radius of the sphere to be very large so that the relaxation time is large. In that case the near horizon geometry of a local patch of the D3-branes will be $AdS_{5}\times S^{5}$ with five form flux[9] and we can study properties of this background by sending observers close to the horizon. [^5] If we scale the 4-form field strength according to (2.5) then the three-brane charge will grow as $\lambda^{4}$[1] and we shall get a locally flat space-time. To get $AdS_{5}\times S^{5}$ geometry we need to keep the three brane charge fixed. Of course if we try to approach infinitely close to the horizon then any experiment will take infinite amount of time from the asymptotic observer’s point of view due to the red shift, but once we fix the accuracy with which we want to do the experiment and estimate the time that will be needed from the asymptotic observer’s point of view to perform such an experiment, we can take the size of the 3-brane system to be sufficiently large so that it does not evolve appreciably during the period of the experiment. Similarly by taking large stacks of spherical M2-branes or M5-branes in M-theory we can produce $AdS_{4}\times S^{7}$ or $AdS_{7}\times S^{4}$ geometries[9]. Since both, the ten dimensional type IIB string theory and the eleven dimensional M-theory, can be produced as backgrounds in the interior of an asymptotically flat four dimensional string theory, obtained by compactifying type IIA string theory on $T^{6}$, we see that all the $AdS_{p}\times S^{q}$ backgrounds listed above can be considered as different states of the same underlying theory. In fact, in this case, we could also wrap the branes along the compact cycles instead of taking them to be spherical since from the asymptotic observer’s viewpoint these are particle like states carrying large but finite energy and charges.
Acknowledgement: This work was supported by the ICTS-Infosys Madhava Chair Professorship and the Department of Atomic Energy, Government of India, under project no. RTI4001.
Appendix A Electrically charged black hole solution in type II on $T^{6}$
In this appendix we shall use the results of [6] to describe the non-rotating black hole solutions in type II string theory on $T^{6}$ carrying fundamental string winding and momentum charges. [^6] Even though the solution in [6] was given for heterotic string theory, the result in type II string theory can be obtained easily from there by truncating the 28 dimensional charge vector to 12 dimensional charge vector and the $28\times 28$ matrix valued scalar moduli $M$ to a $12\times 12$ symmetric $SO(6,6)$ matrix. The solution given here is obtained from the one in [6] by setting to zero the rotation parameter $a$. Also $\Phi$ in [6] is our $2\phi_{4}$. The four dimensional string metric $ds_{4}^{2}$, the four dimensional dilaton $\phi_{4}$ the gauge fields $A_{\mu}$, and the moduli of $T^{6}$, described by a $12\times 12$ symmetric $O(6,6)$ matrix $M$, take the form:
$\displaystyle ds_{4}^{2}$ $=$ $-\Delta^{-1}\rho^{2}(\rho^{2}-2m\rho)dt^{2}+\rho^{2}(\rho^{2}-2m\rho)^{-1}d\rho^{2}+\rho^{2}d\Omega_{2}^{2}\,,$
$e^{-2\phi_{4}}$ $=$ $e^{-2\phi_{0}}\,\Delta^{1/2}\,\rho^{-2}\,,$
$M$ $=$ $I_{12}+\pmatrix{Pnn^{T}&Qnp^{T}\cr Qpn^{T}&Ppp^{T}}\,,$
$A_{t}$ $=$ $-{1\over\sqrt{2}}\,m\,\rho\,\Delta^{-1}\,\pmatrix{\sinh\alpha\{\cosh\beta\rho^{2}+m\rho(\cosh\alpha-\cosh\beta)\}\vec{n}\cr\sinh\beta\{\cosh\alpha\rho^{2}+m\rho(\cosh\beta-\cosh\alpha)\}\vec{p}}\,,$ (A.1)
where
$\displaystyle\Delta$$\equiv$ $\rho^{4}+2m\rho^{3}(\cosh\alpha\cosh\beta-1)+m^{2}\rho^{2}(\cosh\alpha-\cosh\beta)^{2}\,,$
$P$$\equiv$ $2\,\Delta^{-1}\,m^{2}\rho^{2}\sinh^{2}\alpha\sinh^{2}\beta\,,$
$Q$$\equiv$ $-2\Delta^{-1}m\rho\sinh\alpha\sinh\beta\{\rho^{2}+m\rho(\cosh\alpha\cosh\beta-1)\}\,.$ (A.2)
[…]
${1\over 2}\pmatrix{R^{2}+R^{-2}&R^{2}-R^{-2}\cr R^{2}-R^{-2}&R^{2}+R^{-2}}\,.$ ) . (A.3)